OpenStudy (anonymous):

2/y+3 + 3/y-4 = 5/y+6

OpenStudy (anonymous):

\[\frac{ 2 }{ y+3 } + \frac{ 3 }{ y-4 } = \frac{ 5 }{ y+6 }\] Are you supposed to solve for y?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

yah solve for y

OpenStudy (ranga):

Multiply each term by: (y + 3)(y - 4)(y + 6) and simplify

OpenStudy (anonymous):

i did that and got the wrong answer

OpenStudy (ranga):

Why don't you show the steps here and we will see if we can spot the error.

OpenStudy (anonymous):

too much typing

OpenStudy (ranga):

Multiply each term by: (y + 3)(y - 4)(y + 6) 2(y-4)(y+6) + 3(y+3)(y+6) = 5(y+3)(y-4) (2y^2 + 4y - 48) + (3y^2 + 27y + 18) = 5y^2 - 5y - 12 5y^2 + 31y - 30 = 5y^2 - 5y - 12 31y + 5y = 30 - 12 36y = 18 y = 1/2

OpenStudy (anonymous):

oh gees thank you so much!

OpenStudy (ranga):

yw

OpenStudy (anonymous):

but the answer is -11/6

OpenStudy (ranga):

Spotted the mistake.

OpenStudy (ranga):

Multiply each term by: (y + 3)(y - 4)(y + 6) 2(y-4)(y+6) + 3(y+3)(y+6) = 5(y+3)(y-4) (2y^2 + 4y - 48) + (3y^2 + 27y + 54) = 5y^2 - 5y - 60 5y^2 + 31y + 6 = 5y^2 - 5y - 60 31y + 6 = -5y - 60 36y = -66 y = -66/36 = -11/6

OpenStudy (anonymous):

thank you that helped a lot!!!

OpenStudy (ranga):

I am on east coast time and I should not be up this late. I am quitting for the day.

OpenStudy (anonymous):

alright thanks again

OpenStudy (anonymous):

Refer to the attachment from Mathematica.